OpenStudy (anonymous):

Find the points of intersection of the graphs of the equations, and check your results analytically. x^2+x^2=5 x-y=1

OpenStudy (anonymous):

I am begging for your help at this point.

OpenStudy (anonymous):

x^2+x^2=5 if this like this (x^2)+(x^2)=5 then =2x^2=5 ....(1) solving for x here, x=-+sqrt(5/2) ans..... equating this to EQ (2) x-y=1 .....(2) -+sqrt(5/2) -y=1 therefore -sqrt(5/2)-1=y ans... and +sqrt(5/2)-1=y ans...